OpenStudy (anonymous):

Q= a-b/2 solve for a

OpenStudy (anonymous):

Q + b/2 = a

OpenStudy (anonymous):

just add b/2 to both sides

OpenStudy (anonymous):

b/2 to q or a?

myininaya (myininaya):

what does your question mean? b/2 to q or a??

OpenStudy (anonymous):

im trying to solve q= a-b/2 solve for a

OpenStudy (anonymous):

Q= a-b/2 Q + b/2= a - b/2 + b/2 Q + b/2= a

OpenStudy (anonymous):

Do you see how I added b/2 to both sides of the equation? The b/2's on the right side cancel each other out.

myininaya (myininaya):

solve for a means to isolate a lamrc did just that

OpenStudy (anonymous):

A= 1/3 (a+b+c) solve for c

myininaya (myininaya):

multiply both sides by 3 3A=a+b+c now subtract a and subtract c on both sides 3A-a-b=c